# TokenSelect: Efficient Long-Context Inference and Length Extrapolation for LLMs via Dynamic Token-Level KV Cache Selection

### v2

@@ Abstract
The rapid advancement of Large Language Models (LLMs) has driven growing demand for processing extended context sequences in contemporary applications. However, this progress faces two major challenges: performance degradation due to sequence lengths out-of-distribution, and excessively long inference times caused by the quadratic computational complexity of attention. These issues hinder the application of LLMs in long-context scenarios. In this paper, we propose Dynamic Token-Level KV Cache Selection (TokenSelect), a training-free method for efficient and accurate long-context inference. TokenSelect builds upon the observation of non-contiguous attention sparsity, using Query-Key dot products to measure per-head KV Cache criticality at token-level. By per-head soft voting mechanism, TokenSelect selectively involves a few critical KV cache tokens in attention calculation without sacrificing accuracy. To further accelerate TokenSelect, we design the Selection Cache based on observations of consecutive Query similarity and implemented efficient dot product kernel, significantly reducing the overhead. A comprehensive evaluation of TokenSelect demonstrates up to 23.84×23.84$\times$23.84 × speedup in attention computation and up to 2.28×2.28$\times$2.28 × acceleration in end-to-end latency, while providing superior performance compared to state-of-the-art long-context inference methods.
@@
Wei Wu1$\dagger$,$\ddagger$, Zhuoshi Pan2$\dagger$,$\ddagger$, Chao Wang1, Liyi Chen1$\ddagger$, Yunchu Bai1, Tianfu Wang1, Kun Fu3, Zheng Wang3,*, Hui Xiong4,5,* 1University of Science and Technology of China, 2Tsinghua University, 3Alibaba Cloud Computing, 4Hong Kong University of Science and Technology (Guangzhou), 5Hong Kong University of Science and Technology
11footnotetext: Corresponding Authors.22footnotetext: Equal Contribution.33footnotetext: Work during internship at Alibaba Cloud Computing.44footnotetext: Code: https://github.com/pzs19/TokenSelect
@@ 1 Introduction
With the rapid development of large language models (LLMs), the number of parameters is no longer the sole factor significantly affecting model performance. The ability to effectively process longer context information has become one of the key metrics for evaluating LLMs’ capabilities. The latest applications such as cross-document understanding Bai et al. (2024), LLM-powered search systems Sharma et al. (2024), and complex reasoning OpenAI have all placed higher demands on the long-context abilities of LLMs. There are two main difficulties in using pre-trained LLMs for long-context inference. On one hand, LLMs are limited by their context length during pre-training (e.g. Llama 3 only has 8192 tokens). Directly inferencing on longer sequences can lead to severe performance degradation due to reasons including sequence lengths out-of-distribution Xiao et al. (2024b); Han et al. (2024). On the other hand, even if LLMs possess sufficiently large context lengths, the quadratic computational complexity of attention with respect to sequence length makes the response time for long-context inference unbearable.
@@
Figure 1: Distribution of tokens participating in attention computation under different sparsity patterns (indicated by blue dots). TokenSelect can more accurately select critical tokens for attention computation.
@@
Previous works have made numerous attempts to address these difficulties. To extend the context length of LLMs, the current common practice is to perform post-training on long texts Team et al. (2024); Yang et al. (2024a); GLM et al. (2024). However, this approach comes with significant computational costs, particularly in two aspects: the synthesis of high-quality long-text data and the training process on extended sequences. To accelerate long-context inference, many studies focus on the sparsity of attention, attempting to reduce the scale of KV Cache involved in computation. The key to this type of method lies in designing sparse patterns for attention, which can be mainly divided into two categories: one uses predefined sparse patterns Wang et al. (2019); Zaheer et al. (2020); Xiao et al. (2024b); Han et al. (2024), while the other estimates the potential importance of KV Cache during the inference process Zhang et al. (2024c); Oren et al. (2024); Xiao et al. (2024a); Tang et al. (2024b); Jiang et al. (2024), attempting to select relevant KV Cache tokens into attention calculations. However, the design of these sparse patterns is often heuristically based on historical criticality or coarse-grained criticality estimation of tokens, making it difficult to ensure that the selected tokens are truly critical, thus resulting in sub-optimal performance, as shown in Fig. 1.
@@
In this paper, we further observe the non-contiguous sparsity of attention, revealing the importance of designing more fine-grained dynamic sparse patterns. To this end, we propose TokenSelect, a training-free approach that utilizes token-level selective sparse attention for efficient long-context inference and length extrapolation. Specifically, for each Query, TokenSelect dynamically calculates token-level per-head criticality for the past KV Cache and selects the $k$ most critical tokens through our head soft vote mechanism, involving them in the attention calculation. This reduces the scale of attention calculation to a constant length familiar to the model, while maintaining almost all of the long-context information, thereby simultaneously addressing the two main difficulties for long-context inference. To reduce the overhead of token selection, TokenSelect manages the KV Cache in token-level pages Zheng et al. (2024) and designs efficient kernel for token selection based on paged KV Cache management through Triton Tillet et al. (2019). Furthermore, based on our observation of high similarity between consecutive queries, we have designed the Selection Cache, which allows consecutive similar queries to share token selection results, thereby reducing the selection frequency while ensuring its effectiveness.
@@
We evaluate the performance and efficiency of TokenSelect on three representative long-context benchmarks using three open-source LLMs. The experimental results demonstrate that our TokenSelect can achieve up to 23.84×23.84$\times$23.84 × speedup in attention computation compared to FlashInfer flashinfer ai , and up to 2.28×2.28$\times$2.28 × acceleration in end-to-end inference latency compared to state-of-the-art long-context inference method Xiao et al. (2024a). Simultaneously, it provides superior performance on three long-text benchmarks. In summary, we make the following contributions:
@@
  • TokenSelect, a training-free method that achieves accurate and efficient long-context inference and length extrapolation, compatible with mainstream LLM serving systems.
  • Comprehensive evaluations of our method, showing up to 23.84×23.84$\times$23.84 × speedup in attention computation and up to 2.28×2.28$\times$2.28 × acceleration in end-to-end latency while exhibiting superior long-context performance.
@@ 2 Related Works
Long-context LLMs
@@
Due to computational complexity constraints, current LLMs based on Transformers often utilize limited context lengths during pre-training Touvron et al. (2023); Grattafiori et al. (2024); Jiang et al. (2023); Yang et al. (2024a); GLM et al. (2024); AI et al. (2024). To extend the long-context capabilities of LLMs, current methods can be broadly categorized into three approaches Huang et al. (2024); Zhou et al. (2024); Zhao et al. (2024): 1) Modifying positional encodings: A widely adopted method is positional interpolation Chen et al. (2023).  Chen et al. first proposed linear scaling of RoPE Su et al. (2024) to map longer positional ranges within the original training window. Subsequent works bloc97 (2023); emozilla (2023) further improved this method using Neural Tangent Kernel (NTK) theory Jacot et al. (2018), achieving longer context windows while maintaining model performance. Methods like YaRN Peng et al. (2024) and Giraffe Pal et al. (2023) optimize interpolation effects by adjusting frequency components or introducing temperature parameters. 2) Long-context post-training: This approach extends the model’s context length through additional training steps on longer documents after pre-training Yang et al. (2024c); Tian et al. (2024). It has been widely adopted by leading LLMs Team et al. (2024); Yang et al. (2024a); GLM et al. (2024) with the support of sequence parallelism techniques Shoeybi et al. (2020); Jacobs et al. (2023); Liu et al. (2024b). 3) Incorporating additional memory modules: Notable examples include Transformer-XL Dai* et al. (2019), Compressive Transformer Rae et al. (2020), RMT Bulatov et al. (2022) and Infini-attention Munkhdalai et al. (2024). Although these methods have expanded the context length of LLMs, long-context inference still faces the challenge of high computational costs.
Efficient Long-context Inference
@@
In state-of-the-art LLMs serving systems Kwon et al. (2023); Huggingface (2024); NVIDIA (2024); Zheng et al. (2024), technologies such as Flash Attention Dao et al. (2022); Dao (2024) and Paged Attention Kwon et al. (2023) have greatly optimized LLMs inference efficiency by improving GPU I/O bottlenecks. However, in long-context inference scenarios, the quadratic computational complexity of attention with respect to sequence length poses new challenges for LLMs inference. Numerous studies focus on the sparsity of attention, selecting partial KV Cache for attention calculations to improve long-context inference efficiency. Sliding window Wang et al. (2019); Zaheer et al. (2020) is one of the most widely used sparse patterns, reducing complexity to linear by executing attention computations within localized windows. Recent works like StreamingLLM Xiao et al. (2024b) and LM-infinite Han et al. (2024) retain the initial tokens of the sequence in addition to sliding windows, effectively maintaining LLMs’ performance when processing long sequences. While these approaches are simple to implement, they cannot retain information from long contexts. Another approach focuses on dynamic KV Cache selection during inference. Methods like H2O Zhang et al. (2024c), TOVA Oren et al. (2024), FastGen Ge et al. (2024), Scissorhands Liu et al. (2023), and SnapKV Li et al. (2024) evaluate token criticality based on historical attention scores, selecting tokens within a limited budget. However, these methods permanently discard parts of the KV Cache, causing information loss from long contexts. To address this, InfLLM Xiao et al. (2024a) introduces Block Memory Units for KV Cache management, retrieving information from long contexts and offloading less-used blocks to CPU. Similarly, QUEST Tang et al. (2024b) proposes query-aware sparsity at page granularity, while MInference Jiang et al. (2024) optimizes long-context inference using three sparse patterns. Apart from considering all attention heads, some other works Ribar et al. (2024); Lee et al. (2024); Tang et al. (2024a) attempt to focus on only a subset of attention heads. Beyond selection, some other research focuses on KV Cache quantization Liu et al. (2024c); Yang et al. (2024b); He et al. (2024); Kang et al. (2024) and merging Liu et al. (2024a); Wan et al. (2024); Zhang et al. (2024b); Wang et al. (2024). While existing methods have shown progress, opportunities for further improvement remain in achieving optimal accuracy and computational efficiency for real-world deployment.
@@ 3 Preliminaries
In this section, we introduce the inference of LLMs and define the Selective Sparse Attention Problem.
3.1 LLMs Inference
Nowadays, mainstream LLMs are primarily based on the Decoder-only Transformer architecture. Each transformer layer includes a multi-head attention (MHA) and a feed-forward network (FFN). The inference process of LLMs can be divided into two stages: the Prefill Stage and the Decode Stage.
$[\mathbf{Q}_{\text{prefill}},\mathbf{K}_{\text{prefill}},\mathbf{V}_{\text{prefill}}]=\mathbf{X}_{\text{prefill}}\cdot\left[\mathbf{W}_{q},\mathbf{W}_{k},\mathbf{W}_{v}\right],$ (1)
$\mathbf{O}_{\text{prefill}}=\text{softmax}\left(\frac{\mathbf{Q}_{\text{prefill}}\cdot{\mathbf{K}_{\text{prefill}}}^{\top}}{\sqrt{d}}\right)\cdot\mathbf{V}_{\text{prefill}},$ (2)
where $\mathbf{W}_{q},\mathbf{W}_{k},\mathbf{W}_{v}$ are linear projections, []delimited-[][$\cdot$][ ⋅ ] represents tensor concatenation operation, and Eq.(2) is also known as Scaled Dot-Product Attention (SDPA). After these computation, $\mathbf{K}_{\text{prefill}}$ and $\mathbf{V}_{\text{prefill}}$ are stored as the KV Cache for current layer $\mathbf{K}_{\text{cache}}$ and $\mathbf{V}_{\text{cache}}$, and $\mathbf{O}_{\text{prefill}}$ is used for subsequent calculations.
3.2 Selective Sparse Attention
As discussed in the Sec. 1, the high attention sparsity in LLMs suggests sparse attention as a promising solution for long-context inference challenges. Sparse attention can keep the number of tokens participating in attention computations at a constant scale, rather than increasing with sequence length. Given that predefined sparse patterns are detrimental to performance, we aim to dynamically select crucial tokens at each step during the inference process. Therefore, we formalize this problem according to the following definition.
@@ 3.2 Selective Sparse Attention
$\min_{S}\left\|\mathbf{O}-\hat{\mathbf{O}}\right\|_{2}^{2}.$ (7)
Existing works on long-context inference can be categorized under the Selective Sparse Attention Problem, with variations in the design of the selection function $\mathcal{S}$. Zaheer et al. (2020); Xiao et al. (2024b) have developed input-independent selection functions $\mathcal{S}()$, while Zhang et al. (2024c); Oren et al. (2024); Li et al. (2024) propose Query-independent functions $\mathcal{S}(\mathbf{K}_{\text{cache}})$ for improved performance. Current state-of-the-art methods Xiao et al. (2024a); Tang et al. (2024b); Jiang et al. (2024) utilize Query-aware selection functions $\mathcal{S}(\mathbf{Q},\mathbf{K}_{\text{cache}})$. However, these approaches typically operate at a block-level, which limits their effectiveness and overall performance.
4 Motivations and Observations
@@
Previous works on long-context inference have demonstrated the sparsity of attention scores in LLMs, particularly when processing long texts. Recent approaches Xiao et al. (2024a); Jiang et al. (2024); Tang et al. (2024b) partition the KV Cache into non-overlapping blocks, estimating block criticality for sparse attention calculations. These methods assume that tokens with higher attention scores tend to be contiguous. However, our further observations reveal that this assumption does not always hold true in practice. As illustrated in Fig. 2a, attention scores are sparsely distributed at the token-level, with critical tokens not necessarily contiguous. This non-contiguity leads to significant omissions in block-level token selection. Fig. 2b demonstrates that finer selection granularity improves recall of critical tokens, motivating us to perform token-level selection. For token-level selection, an intuitive approach would be to directly select the top-$k$ tokens with the highest attention logits. However, observation in Fig. 2c reveals considerable disparity in the $L_{1}$ norm of attention logits across attention heads. As a result, the selection result tends to be dominated by a few heads with disproportionately large attention logits, driving us to design a more robust selection function that maintains the independence of heads.
@@
As sparsity of attention is dynamic Jiang et al. (2024), token selection should be performed for every Query, which inevitably increases the computational overhead of selective sparse attention. Fortunately, we observe that consecutive Queries exhibit high similarity, as shown in Fig. 3a. Intuitively, when two consecutive Queries are highly similar, their dot products with the Keys will also be similar, leading to substantial overlap in the token selection results. Due to space constraints, we provide an informal lemma about this below. The formal version and corresponding proof can be found in the Appendix A.
@@
Figure 4: The illustration of TokenSelect, which involves calculating per-head criticality using the Paged Dot Product Kernel, performing head soft vote to get selected indices, and executing selective sparse attention via the Paged Attention Kernel.

### v4

@@ Abstract
Rapid advances in Large Language Models (LLMs) have spurred demand for processing extended context sequences in contemporary applications. However, this progress faces two challenges: performance degradation due to sequence lengths out-of-distribution, and excessively long inference times caused by the quadratic computational complexity of attention. These issues limit LLMs in long-context scenarios. In this paper, we propose Dynamic Token-Level KV Cache Selection (TokenSelect), a training-free method for efficient and accurate long-context inference. TokenSelect builds upon the observation of non-contiguous attention sparsity, using QK dot products to measure per-head KV Cache criticality at token-level. By per-head soft voting mechanism, TokenSelect selectively involves a few critical KV cache tokens in attention calculation without sacrificing accuracy. To further accelerate TokenSelect, we design the Selection Cache based on observations of consecutive Query similarity and implemented the efficient Paged Dot Product Kernel, significantly reducing the selection overhead. A comprehensive evaluation of TokenSelect demonstrates up to $23.84\times$ speedup in attention computation and up to $2.28\times$ acceleration in end-to-end latency, while providing superior performance compared to state-of-the-art long-context inference methods.
@@
Wei Wu1222Equal Contribution., Zhuoshi Pan2222Equal Contribution., Kun Fu3, Chao Wang1, Liyi Chen4, Yunchu Bai1, Tianfu Wang5, Zheng Wang3, Hui Xiong5,6111Corresponding Author. 1University of Science and Technology of China,  2Tsinghua University, 3Alibaba Cloud Computing,  4Xiaohongshu Inc., 5The Hong Kong University of Science and Technology (Guangzhou), 6The Hong Kong University of Science and Technology urara@mail.ustc.edu.cn, xionghui@ust.hk
44footnotetext: Code: https://github.com/pzs19/TokenSelect
@@ 1 Introduction
With the rapid development of large language models (LLMs), the number of parameters is no longer the sole factor significantly affecting model performance. The ability to effectively process longer context information has become one of the key metrics for evaluating LLMs’ capabilities. Recent advances—such as cross-document understanding Bai et al. (2024), LLM-powered search systems Sharma et al. (2024), complex reasoning OpenAI , and other cutting-edge LLM developments Wu et al. (2025a); Pan et al. (2025); Wu et al. (2025b); Chen et al. (2024)—have all placed higher demands on the long-context capabilities of LLMs. There are two main difficulties in using pre-trained LLMs for long-context inference. On one hand, LLMs are limited by their context length during pre-training (e.g. Llama 3 only has 8192 tokens). Directly inferencing on longer sequences can lead to severe performance degradation due to reasons including sequence lengths out-of-distribution Xiao et al. (2024b); Han et al. (2024). On the other hand, even if LLMs possess sufficiently large context lengths, the quadratic computational complexity of attention with respect to sequence length makes the response time for long-context inference unbearable.
@@
Figure 1: Distribution of tokens participating in attention computation under different sparsity patterns (blue dots. TokenSelect can more accurately select critical tokens (crimson squares) for attention computation.
@@
Previous works have made numerous attempts to address these difficulties. To extend the context length of LLMs, the current common practice is to perform post-training on long texts Yang et al. (2024a). However, this approach entails significant computational costs, motivating a training-free and effective method that is computationally efficient. To accelerate long-context inference, many studies focus on the sparsity of attention, attempting to reduce the scale of KV Cache involved in computation. The key to this type of method lies in designing sparse patterns for attention, which can be mainly divided into two categories: one uses predefined sparse patterns Wang et al. (2019); Zaheer et al. (2020); Xiao et al. (2024b); Han et al. (2024), while the other estimates the potential importance of KV Cache during the inference process Zhang et al. (2024b); Li et al. (2024); Lee et al. (2024); Tang et al. (2024); Jiang et al. (2024); Sun et al. (2025), attempting to select relevant KV Cache tokens into attention calculations. However, the design of these sparse patterns is often heuristically based on historical criticality or coarse-grained criticality estimation of tokens, making it difficult to ensure that the selected tokens are truly critical, thus resulting in sub-optimal performance, as shown in Fig. 1.
@@
In this paper, we further observe the non-contiguous sparsity of attention, revealing the importance of designing more fine-grained dynamic sparse patterns. To this end, we propose TokenSelect, a training-free approach that utilizes token-level selective sparse attention for efficient long-context inference and length extrapolation. Specifically, for each Query, TokenSelect dynamically calculates token-level per-head criticality for the past KV Cache and selects the $k$ most critical tokens through our head soft vote mechanism, involving them in the attention calculation. This reduces the scale of attention calculation to a constant length familiar to the model, while maintaining almost all of the long-context information, thereby simultaneously addressing the two main difficulties for long-context inference. To reduce the overhead of token selection, TokenSelect manages the KV Cache in token-level pages Zheng et al. (2024) and design efficient kernel for token selection based on paged KV Cache management through Triton Tillet et al. (2019). Furthermore, based on our observation of high similarity between consecutive queries, we have designed the Selection Cache, which allows consecutive similar queries to share token selection results, thereby reducing the selection frequency while ensuring its effectiveness.
@@
We evaluate the performance and efficiency of TokenSelect on three representative long-context benchmarks using three open-source LLMs. The experimental results demonstrate that our TokenSelect can achieve up to $23.84\times$ speedup in attention computation compared to FlashInfer flashinfer ai , and up to $2.28\times$ acceleration in end-to-end inference latency compared to state-of-the-art long-context inference method Xiao et al. (2024a). Simultaneously, it provides superior performance on three long-text benchmarks. In summary, we make the following contributions:
@@
  • TokenSelect, a training-free method that achieves accurate and efficient long-context inference and length extrapolation, which is compatible with mainstream LLM serving systems.
  • Comprehensive evaluations of our method, showing up to $23.84\times$ speedup in attention computation and up to $2.28\times$ acceleration in end-to-end latency while exhibiting superior performance.
@@ 2 Related Works
Long-context LLMs.
@@
Due to computational complexity constraints, current LLMs based on Transformers often utilize limited context lengths during pre-training Touvron et al. (2023); Dubey et al. (2024); Jiang et al. (2023); Yang et al. (2024a); GLM et al. (2024); AI et al. (2024). To extend their long-context capabilities, existing methods can be broadly categorized into three approaches Huang et al. (2024); Zhou et al. (2024); Zhao et al. (2023): 1) Modifying positional encodings: A widely adopted method is positional interpolation Chen et al. (2023).  Chen et al. first proposed linear scaling of RoPE Su et al. (2024) to map longer positional ranges within the original training window. Subsequent works bloc97 (2023); emozilla (2023) further improved this method using Neural Tangent Kernel (NTK) theory Jacot et al. (2018), achieving longer context windows while maintaining model performance. Methods like YaRN Peng et al. (2024) and Giraffe Pal et al. (2023) optimize interpolation effects by adjusting frequency components or introducing temperature parameters. 2) Long-context post-training: This approach extends the model’s context length through additional training steps on longer documents after pre-training Yang et al. (2024b); Tian et al. (2025). It has been widely adopted by leading LLMs Team et al. (2024); Yang et al. (2024a); GLM et al. (2024) with the support of sequence parallelism techniques Shoeybi et al. (2020); Jacobs et al. (2024); Liu et al. (2024b). 3) Incorporating additional memory modules: Notable examples include Transformer-XL Dai* et al. (2019), Compressive Transformer Rae et al. (2020), RMT Bulatov et al. (2022) and Infini-attention Munkhdalai et al. (2024). Although these methods have expanded the context length of LLMs, long-context inference still faces the challenge of high computational costs.
Efficient Long-context Inference.
@@
In state-of-the-art LLMs serving systems Kwon et al. (2023); Zheng et al. (2024), technologies such as Flash Attention Dao (2024) and Paged Attention Kwon et al. (2023) have greatly optimized LLMs inference efficiency by improving GPU I/O bottlenecks. However, in long-context inference scenarios, the computational complexity of attention poses new challenges for LLMs inference. Numerous studies focus on the sparsity of attention, selecting partial KV Cache for attention calculations to improve long-context inference efficiency. Sliding window Wang et al. (2019); Zaheer et al. (2020) is one of the most widely used sparse patterns, reducing complexity to linear by executing attention computations within localized windows. Recent works like StreamingLLM Xiao et al. (2024b) and LM-infinite Han et al. (2024) retain the initial tokens of the sequence in addition to sliding windows, effectively maintaining LLMs’ performance when processing long sequences. While these approaches are simple to implement, they cannot retain information from long contexts. Another approach focuses on KV Cache eviction during inference. Methods like H2Zhang et al. (2024b), TOVA Oren et al. (2024) and SnapKV Li et al. (2024) evaluate token criticality based on historical attention scores, selecting tokens within a limited budget. However, these methods permanently discard parts of the KV Cache, causing information loss from long contexts. To address this, InfLLM Xiao et al. (2024a) introduces Block Memory Units for KV Cache management, retrieving information from long contexts and offloading less-used blocks to CPU. Similarly, QUEST Tang et al. (2024) proposes query-aware sparsity at page granularity, while MInference Jiang et al. (2024) optimizes long-context inference using three sparse patterns. Apart from considering all attention heads, some other works Ribar et al. (2024); Lee et al. (2024) attempt to focus on only a subset of attention heads. While existing methods have shown progress, opportunities for further improvement remain in achieving optimal accuracy and computational efficiency for real-world deployment.
@@ 3 Preliminaries
As discussed in the Sec. 1, the high attention sparsity in LLMs suggests sparse attention as a promising solution for long-context inference challenges, which can keep the number of tokens participating in attention computations at a constant scale. Given that predefined sparse patterns are detrimental to performance, we aim to dynamically select crucial tokens at each step during the inference process. Accordingly, based on the overview of LLM inference presented in Appendix D, we formalize the Selective Sparse Attention Problem as follows.
@@
$\min_{S}\left\|\mathbf{O}-\hat{\mathbf{O}}\right\|_{2}^{2}.\vskip-2.0pt$ (4)
Existing works on long-context inference can be categorized under the Selective Sparse Attention Problem, with variations in the design of the selection function $\mathcal{S}$. Big-Bird and StreamLLM have developed input-independent selection functions $\mathcal{S}()$, while H2O, TOVA and SnapKV propose Query-independent functions $\mathcal{S}(\mathbf{K}_{\text{cache}})$ for improved performance. Current state-of-the-art methods InfLLM, QUEST and MInference utilize Query-aware selection functions $\mathcal{S}(\mathbf{Q},\mathbf{K}_{\text{cache}})$. However, these approaches typically select at a block-level, which limits their effectiveness.
@@
Figure 4: Execution flow of TokenSelect: 1) calculate per‐head criticality via Paged Dot Product Kernel; 2) perform head soft vote to obtain selected indices; 3) execute selective sparse attention via Paged Attention Kernel.
4 Motivations and Observations
@@
Previous works on long-context inference have demonstrated the sparsity of attention scores in LLMs, particularly when processing long texts. Recent approaches (e.g., InfLLM, QUEST and MInference) partition the KV Cache into non-overlapping blocks, estimating block criticality for sparse attention calculations. These methods assume that critical tokens tend to be contiguous. However, our further observations reveal that this assumption does not always hold true in practice. As illustrated in Fig. 2a, attention scores are sparsely distributed at the token-level. This non-contiguity leads to significant omissions in block-level token selection. Fig. 2b demonstrates that finer selection granularity improves recall of critical tokens, motivating us to perform token-level selection. For token-level selection, an intuitive approach would be to directly select the top-$k$ tokens with largest attention logits. However, Fig. 2c reveals considerable disparity in the $L_{1}$ norm of attention logits across attention heads. As a result, the selection result tends to be dominated by a few heads with disproportionately large attention logits, driving us to design a more robust selection function that maintains the independence of heads.
@@
As sparsity of attention is dynamic Jiang et al. (2024), token selection should be performed for every Query, which inevitably increases the computational overhead of selective sparse attention. Fortunately, we observe that consecutive Queries exhibit high similarity, as shown in Fig. 3a. Intuitively, when two consecutive Queries are highly similar, their dot products with the Keys will also be similar, leading to substantial overlap in the token selection results. Due to space constraints, we provide an informal lemma about this below. The formal version and corresponding proof can be found in the Appendix C.